OpenStudy (anonymous):

Which type of probability tool would be an appropriate choice for a simulation that models 4 equally likely outcomes? Standard Die Coin A bag of red and blue marbles Standard deck of cards

OpenStudy (anonymous):

Four suits in the deck of cards sounds attractive. Other choices give p=1/6 or =1/2 or just two outcome types.